# SupertonicTTS: Towards Highly Scalable and Efficient Text-to-Speech System

### v1

@@ Abstract
We present a novel text-to-speech (TTS) system, namely SupertonicTTS, for improved scalability and efficiency in speech synthesis. SupertonicTTS is comprised of three components: a speech autoencoder for continuous latent representation, a text-to-latent module leveraging flow-matching for text-to-latent mapping, and an utterance-level duration predictor. To enable a lightweight architecture, we employ a low-dimensional latent space, temporal compression of latents, and ConvNeXt blocks. We further simplify the TTS pipeline by operating directly on raw character-level text and employing cross-attention for text-speech alignment, thus eliminating the need for grapheme-to-phoneme (G2P) modules and external aligners. In addition, we introduce context-sharing batch expansion that accelerates loss convergence and stabilizes text-speech alignment. Experimental results demonstrate that SupertonicTTS achieves competitive performance while significantly reducing architectural complexity and computational overhead compared to contemporary TTS models. Audio samples demonstrating the capabilities of SupertonicTTS are available at: https://supertonictts.github.io/.
@@ 1 Introduction
To achieve these breakthroughs, researchers have explored diverse approaches to TTS modeling. One prominent direction involves using signal processing features, such as mel spectrograms, as intermediate representations [19, 23, 29, 10]. This approach typically employs two separate models: one to convert text into signal processing features and another to generate a waveform from these features. Another common approach is to leverage discrete tokens derived from neural codec models as intermediate representations [22, 25, 51]. The discrete nature of these features enables language modeling, resulting in improved naturalness, intelligibility, and speaker similarity. Furthermore, disentangled latent spaces are often used to provide control over speech features such as pitch, content, and timbre [21, 8].
@@
Despite significant technical achievements, however, most contemporary TTS systems still require a large number of parameters, massive speech-text training data, and a complex pipeline that includes a grapheme-to-phoneme (G2P) module, a text-speech aligner, or pretrained models for extracting textual and speaker features, as outlined in Table 1. These factors collectively contribute to increased computational overhead during training and inference, and introduce complex interdependencies among system components. Given these challenges, a promising research direction in TTS is to develop a more streamlined pipeline that reduces architectural complexity and training overhead while maintaining competitive performance. The approach of this paper seeks to improve both efficiency and scalability, in line with recent trends in large language models (LLMs) that focus on cost-effective training and optimized inference [33, 9].
@@
We introduce several techniques to significantly enhance scalability and efficiency. First, we design the latent space with a remarkably low dimensionality and compress the latents along the temporal axis before passing them to the text-to-latent module. This strategy significantly reduces the computational overhead of subsequent processes. Second, we introduce context-sharing batch expansion to accelerate loss convergence, offering benefits similar to increasing batch size, but with comparatively lower computational cost. We provide empirical evidence that the proposed batch expansion effectively facilitates accurate text-speech alignment learning. Third, we employ ConvNeXt blocks [34, 47, 39] extensively across all modules to ensure a lightweight and efficient architecture. In addition to these primary contributions, we simplify the TTS pipeline by employing cross-attention mechanisms for text-speech alignment, similar to [31, 37], and by using raw character-level text as input. This design choice removes the need for an external aligner and a grapheme-to-phoneme (G2P) model, which can often become bottlenecks when scaling to new data domains or languages. Furthermore, we refrain from incorporating external pretrained models, thereby reducing architectural dependencies and complexity. Table 1 provides a comprehensive comparison of the proposed system against other TTS models.
@@
Through extensive experimentation, we rigorously evaluate SupertonicTTS and confirm its key contributions. First, we demonstrate its ability to reconstruct high-fidelity speech from compressed latent representations with remarkably fast inference. Second, we show that context-sharing batch expansion accelerates both loss convergence and text-speech alignment learning, offering memory and computational efficiency. Finally, we demonstrate that SupertonicTTS achieves competitive zero-shot TTS performance with just 44 million parameters and extremely fast generation. Audio samples are available at https://supertonictts.github.io/.
@@ 2 Method
SupertonicTTS is built around latent diffusion, which has shown state-of-the-art performance in different generative tasks [45, 36, 43]. More specifically, the training of SupertonicTTS is divided into three phases: first, a speech autoencoder is trained to map input audio into a low-dimensional latent space and reconstruct the original audio. Next, a text-to-latent module learns to generate latent representations that accurately reflect the speech characteristics of both input text and reference speech. Finally, a duration predictor is optimized to estimate the total speech duration based on input text and reference speech.
To ensure fast inference and architectural efficiency, SupertonicTTS primarily employs ConvNeXt blocks [34], which have recently shown competitive performance in speech generation [39, 47, 52]. The entire pipeline is further streamlined by excluding external pretrained models, text-speech aligners, and G2P modules. The overall architecture of SupertonicTTS is depicted in Figure 1 and detailed descriptions of each component are provided in the following sections.
@@ 2.2.2 Context-sharing batch expansion for efficient training
We introduce context-sharing batch expansion to improve the training efficiency of conditional generative models that are based on diffusion or flow-matching algorithms [14, 48, 32]. In conventional training setups, these generative algorithms sample noise and a diffusion timestep to perturb an input variable (forward process) and optimize the models to denoise it using the provided conditional variables (reverse process). In contrast, the proposed batch expansion samples multiple noise instances to produce several perturbed input variables at different diffusion timesteps. Meanwhile, the corresponding conditional variables are reused by duplication 111We assume that the conditional variables are already encoded by individual neural networks.. This strategy mimics the effect of increasing batch size but is computationally more efficient because conditional variables are shared across multiple perturbed inputs (i.e., context-sharing). A pseudo-algorithm is provided in Algorithm 1 for further clarification.
@@ 3.2 Optimization
The text-to-latent module was optimized using the AdamW optimizer for 700k iterations with a batch size of 64 and an expansion factor . The learning rate was initially set to and reduced by half every 300k iterations. Training was conducted on four RTX 4090 GPUs. Latents were normalized using precomputed channel-wise mean and variance statistics before being input into the text-to-latent module. The temporal compression factor was set to 6, resulting in a compressed latent dimension of 144. We set and . During training, reference speech segments were obtained by randomly cropping the input audio, with durations ranging from to . We ensured that the cropped length did not exceed half of the original speech duration.
@@ 4.1 Speech reconstruction
For evaluation of speech reconstruction, we compare the speech autoencoder with BigVGAN222We use the official BigVGAN-v2 checkpoint, pretrained on multilingual speech, environmental sounds, and musical instruments. [10], a state-of-the-art vocoder, using NISQA [38], UTMOSv2 [2], and CREPE [26]. Both NISQA and UTMOSv2 are non-intrusive mean opinion score (MOS) prediction systems to evaluate the perceptual quality of speech signals. Meanwhile, we use CREPE to analyze pitch in speech and compute the F1 score for voiced/unvoiced classification (V/UV F1), which serves as a key metric for identifying artifacts commonly found in non-autoregressive GAN vocoders. Furthermore, we measured the real-time factor (RTF) on an RTX 4090 GPU to quantify inference speed. Our evaluation is conducted using audio samples sourced from the test-clean and test-other subsets of LibriTTS333In this study, all evaluations were conducted using audio samples ranging from 4 to 10 seconds in length..
We report the results in Table 2. In terms of NISQA scores, the speech autoencoder performs competitively, achieving 4.06 on test-clean and outperforming BigVGAN on test-other with a score of 3.76. Similarly, in terms of UTMOSv2 scores, the speech autoencoder slightly lags behind BigVGAN on test-clean but surpasses it on test-other. This indicates that the speech autoencoder can synthesize perceptually high-quality speech. For the F1 evaluation of V/UV classification, BigVGAN achieves the highest scores (0.9735 on test-clean and 0.9620 on test-other), while the speech autoencoder follows closely with 0.9587 and 0.9450, respectively. This suggests that our model introduces minor artifacts affecting voiced/unvoiced classification. Nonetheless, the overall performance remains competitive considering the V/UV F1 scores of other neural vocoders reported in the baseline paper 444The authors of BigVGAN [10] reported that HiFi-GAN [29] and WaveFlow [42] achieved V/UV F1 scores of 0.9300 and 0.9410, respectively, on the dev subsets of LibriTTS.. Notably, the speech autoencoder achieves an inference speed more than 20 times faster than BigVGAN. These results collectively confirm that the speech autoencoder, despite its bottlenecked architecture and low-dimensional latent space, can generate high-quality speech while offering significantly faster inference.
@@ 4.2 Trade-off between the number of function evaluations and synthesis quality
SupertonicTTS synthesizes speech through Euler’s method during inference. By adjusting the number of function evaluations (NFE), we can balance the trade-off between synthesis speed and quality. To quantify this relationship, we generated five samples per utterance from the test-clean subset of LibriSpeech [40] with varying NFE values and evaluated RTF, word error rate (WER), NISQA score [38], and speaker similarity (SIM). SIM was calculated as the cosine similarity between speaker embeddings from the generated speech and the corresponding 3-second reference, using the WavLM-TDNN model [7].
@@ 4.3 Evaluation of context-sharing batch expansion
To assess the impact of the context-sharing batch expansion on training efficiency, we trained four additional SupertonicTTS models using the following batch size () and expansion factor () combinations: (64, 1), (128, 1), (256, 1), and (64, 2). All other architectural and optimization settings remained unchanged. Performance was assessed using two metrics: validation loss and automatic speech recognition (ASR) results. Validation loss was computed on the test-clean subset of LibriTTS by averaging flow-matching losses (Eq. 1) across five timesteps: 0.1, 0.3, 0.5, 0.7, and 0.9. To ensure consistency, the reference mask was kept identical across all evaluations. For evaluation on ASR results, we generated five samples per utterance from the test-clean subset of LibriSpeech, using a 3-second reference speech segment cropped from a different utterance of the same speaker. For inference in the text-to-latent module, we used Euler’s method with 32 steps and a CFG coefficient of 3. The generated samples were transcribed using the CTC-based HuBERT-Large model 555https://huggingface.co/facebook/hubert-large-ls960-ft. [15] and WER were calculated by comparing ground-truth transcriptions with ASR outputs. All transcriptions were normalized using NVIDIA’s NeMo-text-processing [57, 4] before measuring WER.
We present validation loss and WER curves throughout training up to 300k iterations in Figure 2. It can be observed that the validation loss converges faster as the expansion factor increases. This trend is very similar to the effect of increasing the batch size . Interestingly, we found that WER improves more quickly with an increase in the expansion factor, while increasing the batch size leads to degraded WER performance. We hypothesize that this is because the proposed method alleviates the speech-text alignment issue by using the same text and speech pair with different noise and time samples, and simply increasing the number of speech-text pairs does not effectively improve alignment learning. Evaluation results throughout the entire training process are provided in Appendix C.1.
@@ 4.4 Comparison with other zero-shot TTS models
| Ours | 2.64 | 0.83 | 0.472 / 0.915† | 0.23s (RTX 4090) | 915 | 44M |
In this section, we benchmark our proposed model with state-of-the-art TTS systems. Specifically, we evaluate our model against five strong baselines: VALL-E [51], VoiceBox [30], Mega-TTS 2 [20], CLaM-TTS [25], and DiTTo-TTS [31]. These models exhibit impressive TTS performance, even in zero-shot scenarios. For the evaluation of SupertonicTTS, we followed the same experimental setup as described in Section 4.3: five samples were generated per utterance from the test-clean subset of LibriSpeech, using a 3-second reference speech from another utterance of the same speaker, with 32 inference steps and a CFG coefficient of 3. Additionally, we computed character error rates (CER) to assess the finer details of pronunciation in the synthesized speech. To compare speaker similarity with Mega-TTS 2, we also employed WavLM Base+666https://huggingface.co/microsoft/wavlm-base-plus-sv. used in the Mega-TTS 2 paper. Inference time was measured by averaging the processing time for 100 generations of 10-second speech, performed on an RTX 4090.
Table 5 presents the evaluation results, with baseline scores sourced from their publications. In terms of WER, SupertonicTTS performed competitively with a score of 2.64, surpassing VALL-E (5.9) and CLaM-TTS (5.11). For CER, SupertonicTTS achieved a strong result of 0.83, outperforming DiTTo-TTS (0.89). These results suggest that SupertonicTTS can generate speech with minimal skipping, repetition, or mispronunciation issues. Regarding speaker similarity, as measured by WavLM-TDNN, SupertonicTTS scored 0.472 which is lower than VoiceBox (0.662), CLaM-TTS (0.495), and DiTTo-TTS (0.627). This difference may stem from SupertonicTTS’s lack of use of text information for reference speech unlike the other models, potentially affecting zero-shot performance. However, SupertonicTTS achieved a score of 0.915 when evaluated with WavLM Base+, surpassing Mega-TTS 2, which also does not use text for reference speech. Notably, these results were achieved with significantly fewer parameters and less than 2% of TTS dataset size compared to the baselines. Lastly, SupertonicTTS exhibited exceptional inference speed, requiring just 0.23 seconds to generate a 10-second speech sample. This result is considerably faster than the next fastest model, DiTTo-TTS, which requires 1.62 seconds. Overall, SupertonicTTS demonstrates competitive performance with remarkable efficiency in both architecture and training.

### v2

@@ Abstract
We present a novel text-to-speech (TTS) system, namely SupertonicTTS, for improved scalability and efficiency in speech synthesis. SupertonicTTS comprises three components: a speech autoencoder for continuous latent representation, a text-to-latent module leveraging flow-matching for text-to-latent mapping, and an utterance-level duration predictor. To enable a lightweight architecture, we employ a low-dimensional latent space, temporal compression of latents, and ConvNeXt blocks. We further simplify the TTS pipeline by operating directly on raw character-level text and employing cross-attention for text-speech alignment, thus eliminating the need for grapheme-to-phoneme (G2P) modules and external aligners. In addition, we introduce context-sharing batch expansion that accelerates loss convergence and stabilizes text-speech alignment. Experimental results demonstrate that SupertonicTTS achieves competitive performance while significantly reducing architectural complexity and computational overhead compared to contemporary TTS models. Audio samples demonstrating the capabilities of SupertonicTTS are available at: https://supertonictts.github.io/.
@@ 1 Introduction
Despite significant technical achievements, however, most contemporary TTS systems still require a large number of parameters, massive speech-text training data, and a complex pipeline that includes a grapheme-to-phoneme (G2P) module, a text-speech aligner, or pretrained models for extracting textual and speaker features, as outlined in Table 1. These factors collectively contribute to increased computational overhead during training and inference, and introduce complex interdependencies among system components. Given these challenges, a promising research direction in TTS is to develop a more streamlined pipeline that reduces architectural complexity and training overhead while maintaining competitive performance.
@@
We introduce several techniques to significantly enhance scalability and efficiency. First, we design the latent space with a remarkably low dimensionality and compress the latents along the temporal axis before passing them to the text-to-latent module. This strategy significantly reduces the computational overhead of subsequent processes. Second, we introduce context-sharing batch expansion to accelerate loss convergence, offering benefits similar to increasing batch size but with lower computational cost. We provide empirical evidence that the proposed batch expansion effectively facilitates text-speech alignment learning. Third, we employ ConvNeXt blocks [35, 51, 41] extensively across all modules to ensure a lightweight and efficient architecture. In addition to these primary contributions, we simplify the TTS pipeline by employing cross-attention mechanisms for text-speech alignment, similar to [33, 38], and by using raw characters as input. This design choice removes the need for an external aligner and a grapheme-to-phoneme (G2P) model, which can often become bottlenecks when scaling to new data domains or languages. Furthermore, we refrain from incorporating external pretrained models, thereby reducing architectural dependencies and complexity.
@@
Through extensive experimentation, we rigorously evaluate SupertonicTTS and confirm its key contributions. First, we demonstrate its ability to reconstruct high-fidelity speech from compressed latent representations with remarkably fast inference. Second, we show that context-sharing batch expansion accelerates both loss convergence and text-speech alignment learning, offering memory and computational efficiency. Finally, we demonstrate that SupertonicTTS achieves competitive zero-shot TTS performance with just 44 million parameters and extremely fast generation capabilities. Audio samples are available on our demo page.111 https://supertonictts.github.io/.
2 Related work
Modern TTS systems have been developed through various approaches. One major direction utilizes signal processing features, such as mel spectrograms, as an intermediate representation [20, 24, 30, 11, 32]. This approach allows for modular system design where an acoustic model converts text into these features and a vocoder synthesizes a waveform from them. While this simplifies implementation, reliance on hand-crafted features limits latent space exploitation and constrains the model’s representational capacity. Another common approach uses discrete tokens from neural audio codec models as intermediate representations [23, 26, 55]. By leveraging language modeling techniques, this method improves the naturalness, intelligibility, and speaker similarity of synthesized speech. However, errors from the vector-quantization step can degrade speech quality, which is especially critical at low bit-rates. Additionally, the use of residual vector quantization (RVQ) [9, 61] often adds architectural complexity by requiring a prediction of multiple tokens per frame. A third approach focuses on disentangled latent spaces for fine-grained control over diverse attributes of the generated speech [22, 8, 46]. This involves encoding speech into distinct features such as linguistic content, speaker identity, and prosody, and training TTS models to estimate these disentangled features. However, achieving effective disentanglement often requires intricate loss objectives, integrating pretrained models, and a large number of parameters. These requirements can increase engineering efforts and inference latency. In pursuit of efficient and simplified TTS, recent works have sought to avoid complex components like G2P modules, phoneme-level duration modeling, and explicit text-speech aligners [38, 10, 59, 33]. Our work extends these efforts towards enhanced scalability and efficiency by directly processing character-level text, leveraging cross-attention for alignment, and employing a lightweight architecture, without relying on external pretrained models.
@@ 3 Method
SupertonicTTS is built around latent diffusion, which has shown state-of-the-art performance in different generative tasks [49, 37, 45]. More specifically, the training of SupertonicTTS is divided into three phases: first, a speech autoencoder is trained to map input audio into a low-dimensional latent space and reconstruct the original audio. Next, a text-to-latent module learns to generate latent representations that accurately reflect the speech characteristics of both input text and reference speech. Finally, a duration predictor is optimized to estimate the total speech duration based on input text and reference speech. To ensure fast inference and architectural efficiency, SupertonicTTS primarily employs ConvNeXt blocks [35], which have recently shown competitive performance in speech generation [41, 51, 56]. The entire pipeline is further streamlined by excluding external pretrained models, text-speech aligners, and G2P modules. The overall architecture of SupertonicTTS is depicted in Figure 1 and detailed descriptions are provided in the following sections.
@@ 3.2.2 Context-sharing batch expansion for efficient training
We introduce context-sharing batch expansion to improve the training efficiency of conditional generative models that are based on diffusion or flow-matching algorithms [15, 52, 34]. In conventional training setups, these generative algorithms sample noise and a diffusion timestep to perturb an input variable (forward process) and optimize the models to denoise it using the provided conditional variables (reverse process). In contrast, the proposed batch expansion samples multiple noise instances to produce several perturbed input variables at different diffusion timesteps. Meanwhile, the corresponding conditional variables are reused by duplication.222We assume that the conditional variables are already encoded by individual neural networks. This strategy mimics the effect of increasing batch size but is computationally more efficient because conditional variables are shared across multiple perturbed inputs (i.e., context-sharing). A pseudo-algorithm is provided in Appendix C for further clarification.
@@ 4.2 Optimization
The text-to-latent module was optimized using the AdamW optimizer for 700k iterations with a batch size of 64 and an expansion factor . The learning rate was initially set to and halved every 300k iterations. Training was conducted on four RTX 4090 GPUs. Latents were normalized using precomputed channel-wise mean and variance statistics before being input into the text-to-latent module. The temporal compression factor was set to 6, resulting in a compressed latent dimension of 144. We set and . During training, reference speech segments were obtained by randomly cropping the input audio, with durations ranging from to . We ensured that the cropped length did not exceed half of the original speech duration.
@@ 5.1 Speech reconstruction
For evaluation of speech reconstruction, we compare the speech autoencoder with BigVGAN333We use the official BigVGAN-v2 checkpoint, pretrained on multilingual speech, environmental sounds, and musical instruments. [11], a state-of-the-art vocoder, using NISQA [40], UTMOSv2 [2], and CREPE [27]. Both NISQA and UTMOSv2 are non-intrusive mean opinion score (MOS) prediction systems to evaluate the perceptual quality of speech signals. Meanwhile, we use CREPE to analyze pitch in speech and compute the F1 score for voiced/unvoiced classification (V/UV F1), which serves as a key metric for identifying artifacts commonly found in non-autoregressive GAN vocoders. Furthermore, we measured the real-time factor (RTF) on an RTX 4090 GPU to quantify inference speed. Our evaluation is conducted using audio samples sourced from the test-clean and test-other subsets of LibriTTS444In this study, all evaluations were conducted using audio samples ranging from 4 to 10 seconds in length..
We report the results in Table 2. In terms of NISQA scores, the speech autoencoder performs competitively, achieving 4.06 on test-clean and outperforming BigVGAN on test-other with a score of 3.76. Similarly, in terms of UTMOSv2 scores, the speech autoencoder slightly lags behind BigVGAN on test-clean but surpasses it on test-other. This indicates that the speech autoencoder can synthesize perceptually high-quality speech. For the F1 evaluation of V/UV classification, BigVGAN achieves the highest scores (0.9735 on test-clean and 0.9620 on test-other), while the speech autoencoder follows closely with 0.9587 and 0.9450, respectively. This suggests that our model introduces minor artifacts affecting voiced/unvoiced classification. Nonetheless, the overall performance remains competitive considering the V/UV F1 scores of other neural vocoders reported in the baseline paper.555The authors of BigVGAN [11] reported that HiFi-GAN [30] and WaveFlow [44] achieved V/UV F1 scores of 0.9300 and 0.9410, respectively, on the dev subsets of LibriTTS. Notably, the speech autoencoder achieves an inference speed more than 20 times faster than BigVGAN. These results collectively confirm that the speech autoencoder, despite its bottlenecked architecture and low-dimensional latent space, can generate high-quality speech while offering significantly faster inference.
@@ 5.2 Trade-off between the number of function evaluations and synthesis quality
SupertonicTTS synthesizes speech through Euler’s method during inference. By adjusting the number of function evaluations (NFE), we can balance the trade-off between synthesis speed and quality. To quantify this relationship, we generated five samples per utterance from the test-clean subset of LibriSpeech [42] by varying NFE values while keeping a CFG coefficient to 3. These samples were then evaluated in terms of RTF, word error rate (WER), NISQA score [40], and speaker similarity (SIM). SIM was calculated as the cosine similarity between speaker embeddings from the generated speech and the corresponding 3-second reference, using the WavLM-TDNN model [7].
@@ 5.3 Evaluation of context-sharing batch expansion
To assess the impact of the context-sharing batch expansion on training efficiency, we trained four additional SupertonicTTS models using the following batch size () and expansion factor () combinations: (64, 1), (128, 1), (256, 1), and (64, 2). All other architectural and optimization settings remained unchanged. Performance was assessed using two metrics: validation loss and automatic speech recognition (ASR) results. Validation loss was computed on the test-clean subset of LibriTTS by averaging flow-matching losses (Eq. 1) across five timesteps: 0.1, 0.3, 0.5, 0.7, and 0.9. To ensure consistency, the reference mask was kept identical across all evaluations. For evaluation on ASR results, we generated five samples per utterance from the test-clean subset of LibriSpeech, using a 3-second reference speech segment cropped from a different utterance of the same speaker. To address training–inference mismatch due to the absence of punctuation in LibriSpeech, we utilized LibriSpeech-PC [39] for model input. If a corresponding transcription was not available in LibriSpeech-PC, the original LibriSpeech text was capitalized and appended with a period. The generated samples were transcribed using the CTC-based HuBERT-Large model666https://huggingface.co/facebook/hubert-large-ls960-ft. [16] and WER was calculated by comparing the ASR outputs to the original LibriSpeech transcriptions. All transcriptions were normalized using NVIDIA’s NeMo-text-processing [63, 4] before measuring WER.
We present validation loss and WER curves throughout training up to 300k iterations in Figure 2. It can be observed that the validation loss converges faster as the expansion factor increases. This trend is very similar to the effect of increasing the batch size . Interestingly, we found that WER improves more quickly with an increase in the expansion factor, while increasing the batch size leads to degraded WER performance. We hypothesize that this is because the proposed method alleviates the speech-text alignment issue by using the same text and speech pair with different noise and time samples, and simply increasing the number of speech-text pairs does not effectively improve alignment learning. Evaluation results throughout the entire training process are provided in Appendix D.1.
@@ 5.4 Comparison with other zero-shot TTS models
| Ours | 2.64 | 0.83 | 0.472 / 0.915† | 0.23s (RTX 4090) | 945 | 44M |
In this section, we benchmark our proposed model with state-of-the-art TTS systems. Specifically, we evaluate our model against five strong baselines: VALL-E [55], VoiceBox [32], Mega-TTS 2 [21], CLaM-TTS [26], and DiTTo-TTS [33]. These models exhibit impressive TTS performance, even in zero-shot scenarios. For the evaluation of SupertonicTTS, we followed the same experimental setup as described in Section 5.3. Additionally, we computed character error rates (CER) to assess the finer details of pronunciation in the synthesized speech. To compare speaker similarity with Mega-TTS 2, we also employed WavLM Base+777https://huggingface.co/microsoft/wavlm-base-plus-sv. used in the Mega-TTS 2 paper. Inference time was measured by averaging the processing time for 100 generations of 10-second speech, performed on an RTX 4090.
Table 5 presents the evaluation results, with baseline scores sourced from their publications. In terms of WER, SupertonicTTS performed competitively with a score of 2.64, surpassing VALL-E (5.9) and CLaM-TTS (5.11). For CER, SupertonicTTS achieved a strong result of 0.83, outperforming DiTTo-TTS (0.89). These results suggest that SupertonicTTS can generate speech with minimal skipping, repetition, or mispronunciation issues. Regarding speaker similarity, as measured by WavLM-TDNN, SupertonicTTS scored 0.472 which is lower than VoiceBox (0.662), CLaM-TTS (0.495), and DiTTo-TTS (0.627). This difference may stem from SupertonicTTS’s lack of use of text information for reference speech unlike the other models, potentially affecting zero-shot performance. However, SupertonicTTS achieved a score of 0.915 when evaluated with WavLM Base+, surpassing Mega-TTS 2, which also does not use text for reference speech. Notably, these results were achieved with significantly fewer parameters and less than 2% of TTS dataset size compared to the baselines. Lastly, SupertonicTTS exhibited exceptional inference speed, requiring just 0.23 seconds to generate a 10-second speech sample. This result is considerably faster than the next fastest model, DiTTo-TTS, which requires 1.62 seconds. Overall, SupertonicTTS demonstrates competitive performance with remarkable efficiency in both architecture and training. A subjective listening test in Appendix D.2 further confirms its perceptual competitiveness in speech naturalness and speaker similarity.